The length of the frame is three times the width and it's diagonal measurement is 160 cm, what are the dimensions of the frame?
Answer by lwsshak3(11628)

The length of the frame is three times the width and it's diagonal measurement is 160 cm, what are the dimensions of the frame?
***
let x=width of rectanglar frame
3x=length of rectangular frame
By the pythagorean theorem:
√(width^2+length^2)=diagonol
√((x^2)+(3x)^2)=√x^2+9x^2=√(10x^2)=160
√(10x^2)=160
√10x=160
x=160/√10≈50.60
3x=3*160/√10≈151.79
width of rectanglar frame≈50.60 cm
length of rectangular frame≈151.79 cm